OpenStudy (anonymous):

How many molecules are there in 15.7 liters of chlorine gas at STP?

OpenStudy (anonymous):

n= V/Vm n=N/L V/Vm=N/L N(Cl2) = (V(Cl2)/22,4 dm^3 mol^-1)*6,022*10^23 mol^-1